# NeuroSim V1.5: Improved Software Backbone for Benchmarking Compute-in-Memory Accelerators with Device and Circuit-level Non-idealities

### v1

@@ IV-A Design Space Exploration
We provide example performance metrics for the default configuration in NeuroSim using the selected RRAM device. As shown in Table III, we achieve 11.6 TOPS at 21.3 TOPS/W while processing 7,770 frames per second for CIFAR-100 images. Through systematic exploration of the design space, we can find designs that significantly outperform this baseline. Results are shown in Fig. 4. From this exploration, we identify several key trends that characterize optimal CIM architectures:
@@
    ADC Precision: The Pareto-optimal designs cluster in the 5--8 bit ADC precision range. Our analysis shows that ADC precision can generally be reduced by 1-bit from the lossless precision without significant accuracy degradation, and by 2-bits in certain configurations.
@@
    Array Dimensions: Optimal designs utilize array sizes between $32\times 32$ and $128\times 128$ cells. Smaller arrays reduce ADC precision requirements but necessitate additional accumulation circuitry, while larger arrays necessitate larger ADCs, incurring costs in latency and energy. The designs that achieve $>$40 TOPS/W used array sizes of either $32\times 32$ or $64\times 64$.
@@
    Memory Cell Precision: 2-bit and 3-bit MLC configurations dominate the Pareto frontier, offering an effective balance between storage density and programming reliability. These configurations achieve the highest energy efficiency (TOPS/W) for any given accuracy target.
The interplay between these parameters creates distinct architectural sweet spots. For example, the 8b/8b/64/6/2 configuration (input/weight/array size/ADC/MLC) achieves 75% accuracy while maintaining high energy efficiency. Designs with reduced precision (6b/6b) can push efficiency above 40 TOPS/W at the cost of $\sim 5\%$ accuracy degradation. The ability to rapidly evaluate these configurations through NeuroSim V1.5 enables the optimization of architectural parameters based on specific device and circuit designs.
Next, we demonstrate the improved noise-modeling capabilities in the new behavioral simulator.
@@ IV-B Noise Modeling Case Studies
Figure 5: Inference accuracy of VGG8 on CIFAR-10, ResNet18 on CIFAR-100, and ResNet50 on ImageNet using V1.5 across different RRAM (single-level cell) variations.
@@ IV-B1 D2D Variation
We evaluated inference accuracy across multiple neural networks using NeuroSim V1.5’s comprehensive noise modeling capabilities, including D2D variation, memory state drifting, stuck-at-faults and circuit-level MAC output noise. To systematically assess accuracy degradation, we selected network architectures of increasing complexity: VGG8 for CIFAR-10, ResNet18 for CIFAR-100, ResNet50 for ImageNet, and Swin Transformer Tiny for ImageNet. Following established methodology [62], we excluded the first convolutional layer from quantization in all simulations.
Figure 5 reveals that accuracy degradation becomes more severe with increasing network complexity and dataset difficulty. VGG8 on CIFAR-10 maintains reasonable accuracy until LRS variation reaches 20%. In contrast, ResNet18 on CIFAR-100 and ResNet50 on ImageNet show significant accuracy drops at just 10% and 5% LRS variation, respectively. Most notably, Swin Transformer Tiny demonstrates the lowest tolerance to device-level variations, highlighting the need for additional research into improving transformer model robustness for analog computing implementations.
Figure 6: Inference accuracy of VGG8 on CIFAR-10 regarding random device conductance drifting, drifting to Gmax, and drifting to Gmin with 1b cells.
@@ IV-B2 Conductance Drift
Fig. 6 examines how three different types of conductance drift (modeled by Equation 5) affect inference accuracy. First, random drift shows better accuracy retention compared to drifting toward minimum or maximum conductance. This improved resilience occurs because cells at minimum conductance cannot drift lower, while cells at maximum conductance cannot drift higher, effectively limiting the total conductance change possible under random drift. Second, drifting toward maximum conductance maintains better accuracy than drifting toward minimum conductance. This asymmetry arises from Equation 5, where the drift coefficient v determines the drift direction. For the same retention time t and drifting speed, drifting toward minimum conductance results in a larger conductance difference compared to drifting toward maximum, leading to a larger deviation of MAC outputs.
Figure 7: Inference accuracy of VGG8 and ResNet18 under different stuck-at-fault rates.
@@ IV-B3 Stuck-at-Faults
NeuroSim V1.5 adds modeling capabilities for stuck-at-fault (SAF) errors, extending beyond the D2D variation and drift modeling present in V1.4. Fig. 7 demonstrates how SAF errors impact inference accuracy across different networks, with SAF rates bounded by realistic RRAM fabrication constraints: 1.75% for LRS and 9.0% for HRS [42]. Our analysis reveals several key findings. First, both 1-bit and 2-bit cells show similar accuracy degradation at equivalent SAF rates, suggesting that cell precision does not significantly influence SAF sensitivity. However, network complexity strongly affects SAF tolerance - ResNet-50 on ImageNet loses accuracy even at minimal SAF rates, while VGG8 and ResNet18 maintain limited functionality only at the lowest SAF rates. Notably, SAF errors cause more severe accuracy degradation compared to other noise sources like D2D variation or conductance drift. This heightened sensitivity to SAF errors aligns with previous research findings [42], highlighting the critical importance of minimizing stuck faults during device fabrication.
Figure 8: (a) Actual MAC output with 2b FeFET current domain CIM [43] under 50mV Vth variation; (b) Actual MAC output with 2b FeFET charge domain CIM [43] under 50mV Vth variation; (c) Actual MAC output of RRAM CIM tape-out [41]; (d) Actual MAC output with 1b nvCap charge domain CIM [33, 18] under 50mV Vth variation; (e) Inference accuracy of various neural network algorithms across CIM macro A-D.
@@ IV-B4 MAC Output Noise
CIM A and B [43] employ 2b FeFET technology with different computing modes: CIM A performs current-mode computation with FeFETs directly, while CIM B implements charge-based computation using FeFET-modulated capacitors. Their noise characteristics were derived from Monte-Carlo SPICE simulations. CIM C [7] provides real silicon validation through direct measurements from a fabricated RRAM-based CIM chip. For these three designs, each MAC output level is characterized by unique mean and standard deviation values. CIM D [33, 18] uses 28 nm nvCap with charge-mode computation, where SPICE simulations revealed thermal noise as the dominant source, here, a uniform variation across output levels was used.
@@
Figs. 8(a-d) illustrate the statistical distribution of actual versus ideal MAC outputs for representative cases. While CIM A, B, and D show tighter error bounds compared to CIM C, all designs achieve similar accuracy on smaller networks. Fig. 8(e) compares inference accuracy across architectures of increasing complexity. VGG8 and ResNet18 maintain reasonable accuracy across all implementations, with CIM C showing slightly higher degradation ( 2%) due to increased output variation. For ResNet50, only CIM B and D - which achieve lower standard deviations in their MAC errors - maintain high accuracy. The Swin Transformer exhibits particular sensitivity to output noise, with only the nvCap-based CIM D achieving software-comparable performance.
@@
These results validate NeuroSim V1.5’s noise modeling framework as a versatile platform for evaluating CIM designs using either simulation data or silicon measurements. By providing a standardized methodology for capturing MAC output statistics, the framework enables direct comparison of different technologies and compute circuits under realistic operating conditions.
IV-C Runtime Evaluation and Simulator Comparison
We first compare NeuroSim V1.5 against existing open-source ACIM simulation frameworks to highlight key capabilities and performance improvements (Table IV). CrossSim [37] provides detailed device-level accuracy simulation but faces significant runtime overhead when modeling noise effects. AIHWKit [63] specializes in noise-aware training but offers limited support for general noise modeling and architectural exploration. While NeuroSim V1.4 [54] includes PPA analysis and has some noise modeling, it was constrained by slower inference speeds and limited network support.
@@ IV-C Runtime Evaluation and Simulator Comparison
Supported network types*
@@
  • * PPA evaluation excluded.
The runtime is primarily influenced by the precision parameters that determine computation complexity through Equation 3: DAC precision affects the number of input cycles, cell bit-precision (MLC) determines the number of weight components, and noise modeling adds statistical sampling overhead. With 8b/8b/None configuration (DAC / MLC / Noise) shown in Table V, V1.5 achieves a baseline runtime of 0.95 ms/image for VGG8 and CIFAR-10. When using 1b cells to represent 8b weights, runtime increases $\sim$4$\times$ to 4.1 ms/image due to additional bit-slicing operations. Adding device noise results in negligible runtime increase and adding per-MAC output noise increases runtime to only 5.1 ms/image.
NeuroSim V1.5 achieves this dramatic improvement in simulation efficiency through three key advances: a PyTorch-optimized ACIM behavioral simulator, noise modeling with pre-characterized statistical noise modeling and TensorRT integration for efficient quantization. As shown in Fig. 9, these optimizations deliver a 6.51$\times$ speedup over V1.4 for CIFAR-10 inference with 1-bit cells.
@@
Tables V and VI further demonstrate V1.5’s efficient noise modeling capabilities on larger networks towards CIFAR-100 and ImageNet datasets. Compared to CrossSim, which shows a significant slowdown when incorporating device noise, V1.5’s statistical approach adds minimal overhead - just 1.3$\times$ for uniform noise and 3.1$\times$ for output-dependent noise simulation. This efficiency stems from our pre-characterized noise distribution approach, eliminating the need for detailed circuit simulation during inference discussed in section III-B.
@@
These advances establish NeuroSim V1.5 as a comprehensive development platform that bridges detailed device/circuit-level analysis with system-level architectural exploration. The framework’s efficiency and adaptability make it particularly valuable as the field moves toward more complex network architectures requiring extensive accuracy and performance optimization.
TABLE VI: Inference Runtime (Seconds/Image) Comparison of ResNet50 on ImageNet (8b activation, 8b weight) between CrossSim, AIHWKit, and NeuroSim V1.5
@@
V Conclusion
The growing complexity of AI models places increasing demands on CIM accelerator design, requiring tools that can efficiently evaluate both accuracy and hardware implementation trade-offs. This work presents NeuroSim V1.5, which combines accuracy analysis with power, performance, and area estimation to support systematic design space exploration of CIM architectures.
The framework’s integration with TensorRT enables evaluation of pre-trained networks without modification, while our flexible noise modeling methodology supports both detailed device-level analysis and efficient MAC-level noise injection, validated against published SPICE simulations and silicon measurements. These capabilities, combined with up to 6.5× faster runtime compared to NeuroSim V1.4, enable more extensive design space exploration than previous versions.
Through our case studies with a variety of device technologies and neural network topologies, we demonstrate how NeuroSim V1.5 can provide insights into CIM design trade-offs:

### v2

@@ III-F4 Runtime Evaluation
IV Results and Case Studies
IV-A Design Space Exploration
We demonstrate NeuroSim V1.5’s design space exploration capability using a 22 nm RRAM [6] platform. Following Table I, we evaluate accuracy-efficiency tradeoffs across architectural choices including array size, precision, and data representation.
@@ IV-A Design Space Exploration
Table II shows example performance metrics for the default configuration: 11.6 TOPS at 21.3 TOPS/W while processing 7,770 frames per second for CIFAR-100 images. Systematic exploration (Fig. 5) reveals designs that significantly outperform this baseline, identifying key trends for optimal CIM architectures:
@@
    ADC Precision: Pareto-optimal designs cluster in 5-8 bit range. ADC precision can be reduced by 1-bit from lossless precision without significant accuracy loss, and by 2-bits in certain configurations.
@@
    Array Dimensions: Optimal designs use array sizes between $32\times 32$ and $128\times 128$ cells. Smaller arrays reduce ADC precision requirements but need more accumulation circuitry, while larger arrays require larger ADCs, incurring latency and energy costs. Designs achieving >>40 TOPS/W used $32\times 32$ or $64\times 64$ arrays.
@@
    Memory Cell Precision: 2-bit and 3-bit MLC configurations dominate the Pareto frontier, balancing storage density and programming reliability while achieving highest energy-efficiency for any given accuracy target.
Across the explored configurations, Pareto-optimal designs converge toward 6-7 bit ADC precision regardless of other parameter choices. This range represents the balance point between ADC overhead (energy, latency, area) and throughput. Multiple parameter combinations achieve similar efficiency at this ADC precision: 1-bit cells with 128 rows, 2-bit cells with 64 rows, or 4-bit cells with 32 rows. The choice between these configurations then shifts to device multi-state capability and rate of variations. Current ADC designs dominate the energy and latency budget in ACIM arrays. If more efficient ADC architectures become available—through techniques such as time-domain sensing, or quantization schemes for reduced-precision —the optimal design point would shift toward larger array sizes or higher-precision MLC devices that are currently constrained by ADC overhead. Next, we demonstrate the improved noise-modeling capabilities in the new behavioral simulator.
Figure 6: Inference accuracy of VGG8 on CIFAR-10, ResNet-18 on CIFAR-100, ResNet-50, and Swin-T on ImageNet with different degrees of 1b RRAM device variation. ”XRow” describes the number of rows activated in parallel.
@@ IV-B Noise Modeling Case Studies
We evaluated inference accuracy across multiple neural networks using NeuroSim V1.5’s noise modeling capabilities, including D2D variation, memory state drifting, stuck-at-faults and circuit-level MAC output noise. We selected network architectures of increasing complexity: VGG8 for CIFAR-10, ResNet-18 for CIFAR-100, ResNet-50 for ImageNet, and Swin Transformer Tiny for ImageNet. Following established methodology [61], we excluded the first convolutional layer from quantization.
@@ IV-B1 D2D Variation
To explore the effect of memory state variation on inference accuracy, we use a 1-bit RRAM baseline design and select several variation rates in the HRS and LRS ranging from 1% to 40%. Based on previous measurement results, we use a higher variation in the HRS than the LRS.
[…]
Figure 7: Inference accuracy of VGG8/CIFAR-10 regarding random device conductance drifting, drifting to Gmax, and drifting to Gmin with 1b cells.
@@ IV-B2 Conductance Drift
Fig. 7 examines how three different types of conductance drift (modeled by Equation 5) affect inference accuracy. Drifting toward maximum conductance (Gmax) maintains the best accuracy retention. Random drifting shows the next best resilience; this occurs because cells at minimum conductance cannot drift lower, while cells at maximum conductance cannot drift higher, effectively limiting the total conductance change. Drifting toward minimum conductance (Gmin) results in the largest accuracy degradation. This asymmetry between Gmax and Gmin drift arises from Equation 5; for the same retention time $t$ and drifting speed $|v|$, drifting toward Gmin can result in a larger absolute conductance change (if starting from a higher state) compared to drifting toward Gmax (if starting from a lower state), leading to a larger deviation of MAC outputs.
@@ IV-B3 Stuck-at-Faults
NeuroSim V1.5 adds stuck-at-fault (SAF) error modeling, extending beyond V1.4’s D2D variation and drift modeling. Fig. 8 shows SAF error impact on inference accuracy across different networks, with SAF rates bounded by realistic RRAM fabrication constraints: 1.75% for LRS and 9.0% for HRS [62]. Both 1-bit and 2-bit cells show similar accuracy degradation at equivalent SAF rates, suggesting cell precision does not significantly influence SAF sensitivity. However, network complexity strongly affects SAF tolerance - ResNet-50 on ImageNet loses accuracy even at minimal SAF rates, while VGG8 and ResNet-18 maintain limited functionality only at the lowest SAF rates. SAF errors cause more severe accuracy degradation than other noise sources like D2D variation or conductance drift. This aligns with previous findings [62], highlighting the critical importance of minimizing stuck faults during fabrication.
Figure 8: Inference accuracy of VGG8 and ResNet-18 under different stuck-at-fault rates (1b RRAM, 128x128 array, 7b ADC).
@@ IV-B4 MAC Output Noise
CIM A and B [34] employ 2b FeFET technology with different computing modes: CIM A performs current-mode computation with FeFETs, while CIM B implements charge-based computation using FeFET-modulated capacitors. Their noise characteristics were derived from Monte-Carlo SPICE simulations. CIM C [7] uses silicon measurements from a fabricated RRAM-based CIM chip. For these three designs, each MAC output level is characterized by unique mean and standard deviation values. CIM D [27, 18] uses 28 nm nvCap with charge-mode computation, where SPICE simulations revealed thermal noise as the dominant source, here, a uniform variation across output levels was used.
@@
Figs. 9(a-d) illustrate the statistical distribution of actual versus ideal MAC outputs for representative cases. While CIM A, B, and D show tighter error bounds compared to CIM C, all designs achieve similar accuracy on smaller networks. Fig. 9(e) compares inference accuracy across architectures of increasing complexity. VGG8 and ResNet-18 maintain reasonable accuracy across all implementations, with CIM C showing slightly higher degradation ( 2%) due to increased output variation. For ResNet-50, only CIM B and D - which achieve lower standard deviations in their MAC errors - maintain high accuracy. The Swin Transformer exhibits particular sensitivity to output noise, with only the nvCap-based CIM D achieving software-comparable performance.
@@
These results also validate the circuit expert mode’s statistical modeling approach. By characterizing MAC output distributions from SPICE simulations or silicon measurements and injecting noise accordingly, the framework captures the aggregate impact of device and circuit non-idealities without requiring per-cell physical simulation. By providing a standardized methodology for capturing MAC output statistics, the framework enables direct comparison of different technologies and compute circuits under realistic operating conditions.
Figure 9: (a) Actual MAC output with 2b FeFET current domain CIM [34] under 50mV Vth variation; (b) Actual MAC output with 2b FeFET charge domain CIM [34] under 50mV Vth variation; (c) Actual MAC output of an RRAM tape-out [33]; (d) Actual MAC output with 1b nvCap charge domain CIM [27, 18] under 50mV Vth variation; (e) Inference accuracy of various neural network algorithms across CIM macro A-D.
IV-C Noise Sensitivity Analysis of ViTs
As shown in Fig. 6, Swin-T exhibits lower tolerance to analog noise compared to ResNet-50. To understand the underlying mechanisms, we analyze errors at both the layer and circuit levels using device-to-device (D2D) variations as the primary noise source.
IV-C1 Related Work on Transformer Sensitivity
Prior work attributes transformer sensitivity in analog CIM to non-uniform activation distributions causing quantization challenges [50, 49], heightened input/output interface noise susceptibility [63], and MSB-cycle error amplification in bit-serial computation [64].
IV-C2 Layer-wise Error Analysis
To identify whether specific layers or architectural components drive Swin-T’s higher noise sensitivity, we compute the root mean square error (RMSE) in each layer output under the 4%/2% D2D variation condition. RMSE is computed as $\sqrt{\text{mean}((\hat{y}-y)^{2})}/\sqrt{\text{mean}(y^{2})}$, where $y$ is the ideal floating-point layer output and $\hat{y}$ is the noisy output. Tests were conducted for both ResNet-50 and Swin-T on ImageNet with 1-bit cells, an array size of $\times128$, and 7-bit ADC.
Figure 10: Layer-wise RMSE for ResNet-50 vs Swin-T under 4%/2% (HRS/LRS) D2D variation. Swin-T shows higher error variance with attention layers (triangles) exhibiting greater sensitivity than non-attention layers (squares).
[…]
Figure 11: CIM ADC output distribution under ResNet-50 and Swin-T using 8b input and 8b weight with 128x128 arrays and 4%/2% (HRS/LRS) D2D variations. Collected from 100 images in ImageNet dataset.
IV-C3 ADC Error Analysis
The layer-level RMSE analysis reveals error magnitudes after accumulating partial sums and dequantizing to floating-point. To understand the root cause of these errors, we examine errors at the ADC output stage, where integer MAC partial sums are quantized. We define error rate as the percentage of partial sums where the noisy ADC output deviates from the ideal value, computed per expected output bin. Analysis of ADC output distributions reveals that Swin-T generates significantly higher average ADC outputs compared to ResNet-50 (Fig. 11). Larger ADC outputs experience increased error rates due to D2D variations, as shown in Fig. 12. Consequently, Swin-T accumulates a higher fraction of erroneous ADC outputs across all layers, contributing significantly to accuracy degradation.
The increased error rate at higher ADC output values stems from the accumulation of device variations across activated cells. Each memory cell contributes independent D2D variation to the column current/charge. When more cells are activated (corresponding to larger partial sums), the total variance increases with the number of contributing devices. Swin-T’s activation and weight distributions produce systematically higher partial sums than ResNet-50, placing a larger fraction of computations in the high-error region of Fig. 12.
Prior work has attributed transformer sensitivity in analog CIM primarily to MSB error amplification during bit-serial computation [64]. While MSB errors do have a disproportionate impact on output magnitude, this mechanism alone does not explain why ViTs exhibit lower noise tolerance than CNNs—both architectures use identical bit-serial processing. Our analysis indicates that the distribution of ADC outputs, determined by activation sparsity and weight magnitude statistics, is the distinguishing factor. CNNs with ReLU activations produce sparser outputs concentrated at lower ADC values, while transformers with GELU activations and larger weight magnitudes shift the distribution toward higher values where error rates increase.
Figure 12: Error rate as a function of expected ADC output value for ResNet-50 and Swin-T with 128x128 arrays under 4%/2% (HRS/LRS) D2D variations. Collected from 100 images on the ImageNet dataset.
IV-C4 Mitigation Strategies
To improve Swin-T accuracy under analog noise, we decreased the size of the memory arrays to $32\times 128$. This effectively reduces the ADC outputs to values less than 32, where the rate of errors is below 2% (Fig. 12). As shown in Fig. 6, this modification achieves accuracy close to the software baseline at 4%/2% D2D variation; however, accuracy quickly falls off at higher variation. To achieve accuracy comparable to ResNet-50 under D2D variations, we need to reduce the number of active rows in parallel to 8. The reduction in active rows does not require a new hardware configuration, as each group of 8 rows in an array are activated sequentially. We quantify the tradeoff in performance in (Table III). We find that Swin-T using $32\times 128$ configuration achieves near parity with ResNet-50 in throughput and energy efficiency but with $5.4\times$ worse area efficiency. While comparable accuracy is achievable, CNNs remain preferable for CIM-based vision accelerators.
Alternative approaches to improve transformer resilience to ACIM noise include: (1) hardware-aware training with noise injection [65]; (2) input and weight rescaling methods to shift their distributions to fit within quantization constraints [66]; (3) Training low-rank adapters to correct errors arising from CIM noise [67]. ResNet-50’s inherent advantages—weight distributions closer to zero and ReLU-induced sparsity—naturally produce lower ADC outputs and improved noise tolerance. Applying techniques to encourage these input and weight distributions in transformer architectures can bridge this robustness gap without requiring hardware modifications.
TABLE III: 22 nm RRAM ResNet-50 and Swin-T PPA under different row parallelism
Model Array Read Rows TOPS TOPS / mm2 TOPS/W FPS
ResNet50 128x128 128 4.84 0.0147 5.04 378
Swin-T 32x128 32 4.83 0.00272 4.84 388
Swin-T 32x128 8 3.46 0.00192 4.34 278
IV-D PPA Breakdown of Swin-T
Next, we evaluate the PPA of Swin Transformer [68] on ImageNet. The linear and convolution layers are implemented in 22 nm RRAM-based ACIM while the Q,K,V matrix multiplications in attention are computed in 22 nm SRAM-based DCIM. Both CIM arrays are configured with a size of $32\times 128$. Fig. 13 presents the detailed PPA breakdown. For area, DCIM accumulation including adder trees and shift adders dominates, resulting in $1.5\times$ larger area than ACIM. However, the energy profile differs: the ADC in ACIM contributes the largest portion of energy consumption compared to significantly lower consumption of DCIM computation. This is because the linear and convolutional layers have higher arithmetic intensity than attention, requiring a higher usage of the ACIM arrays and thus higher energy consumption. This asymmetry suggests that for area-constrained designs, minimizing DCIM-computed operations (e.g., offloading some attention to external processors) may be more effective, while for energy-constrained designs, ADC optimization or reduced-precision sensing offers larger gains if accuracy can be maintained via quantization-aware tuning methods. For the critical path latency, the adder tree in DCIM and the ADC in ACIM show comparable delays. The peripheral circuits in ACIM array introduce additional overhead with longer overall latency compared to DCIM. A potential optimization for DCIM-based attention is leveraging finer-grained pipelining at the window level or head level with carefully designed dataflow and higher memory array reuse, which can reduce area while maintaining high-throughput.
Figure 13: PPA breakdown of Swin-T (25M parameters). The majority of area is consumed by DCIM adder trees within memory arrays. Array size was reduced to 32x128 to decrease adder tree size. The majority of energy consumption and latency arises from ACIM ADC.
IV-E Runtime Evaluation and Simulator Comparison
Finally, we compare NeuroSim V1.5 against existing open-source ACIM simulators (Table IV). CrossSim [30] excels at device-level physical simulation with high fidelity but higher runtime, especially for complex noise. AIHWKit [63] focuses on noise-aware training with phase-change memory. NeuroSim V1.4 [54] introduced integrated PPA and some noise modeling but had slower inference and narrower network support. NeuroSim V1.5 distinguishes itself by integrating flexible, gpu-accelerated noise modeling, and vision transformer support, with circuit-level PPA estimation. This enables a wider range of noise, quantization, and circuit parameters to be explored than in other open-source simulators.
@@ IV-E Runtime Evaluation and Simulator Comparison
Supported Network Types
@@
Table V details runtime scaling with precision and noise configuration. NeuroSim V1.5 achieves 0.95 ms/image baseline for VGG8/CIFAR-10, increasing to 4.1 ms/image with 1b cells due to bit-slicing overhead. Device noise adds negligible cost; per-MAC output noise increases runtime to 5.1 ms/image.
[…]
@@
[…]
These advances improve NeuroSim’s functionality as a design space exploration platform across memory devices, computing methods, and technology nodes. Evaluating both network accuracy and performance is essential for optimal CIM hardware design, as ignoring either metric provides an incomplete picture and can exaggerate throughput and energy-efficiency.
TABLE VI: Inference Runtime (Seconds/Image) Comparison of ResNet-50 on ImageNet (8b activation, 8b weight) between CrossSim, AIHWKit, and NeuroSim V1.5
@@
V Conclusion and Future Work
The growing complexity of AI models places increasing demands on CIM accelerator design, requiring tools that can efficiently evaluate both accuracy and hardware performance. This work presents NeuroSim V1.5, which combines accuracy analysis with power, performance, and area estimation. A key contribution is the heterogeneous ACIM/DCIM architecture extending CIM support to transformer models—increasingly important as transformers dominate AI workloads from vision to language.
[…]
Through our case studies with various device technologies and neural network topologies, we demonstrate how NeuroSim V1.5 can provide insights into CIM design trade-offs: